Please help me solve this equation: y = 3.2 - 0.2y
Answer by checkley71(8403)

Y=3.2-0.2Y OR THIS IS BEST SOLVED USING FRACTIONS SO
Y=3 2/10-(2/10)Y OR (12/10)Y=32/10 OR Y=(32/10)/(12/10) OR Y=32/12 OR Y=8/3
PROOF
8/3=32/10-(2/10)(8/3) OR 8/3=32/10-16/30 OR 80/30=96/30-16/30 OR 80/30=80/30
OR 8/3=8/3